THEORETICAL DETERMINATION AND ASSIGNMENT OF THE VIBRATIONAL SPECTRA OF ACTINOCENES: An(η^{8}-C_{8}H_{8})_{2} (An = Th, Pa, U, Np, Pu, Am)

\underline{JUN LI} and BRUCE E. BURSTEN, Department of Chemistry, The Ohio State University, 100 West 18^{th} Avenue, Columbus Ohio 43210.

Relativistic density functional methods have been used to calculate the optimized geometries and vibrational frequencies of actinocenes, An(η^{8}-C_{8}H_{8})_{2} (An = Th, Pa, U, Np, Pu, Am). The influences of scalar and spin-orbit relativistic effects on the geometries and vibrational frequencies of these organoactinide complexes are discussed. The calculated scalar-relativistic vibrational frequencies and absorption intensities are in excellent agreement with experimental results. Clear-cut assignments of all the infrared and Raman vibrational modes of actinocenes are presented.